OpenStudy (anonymous):

A particle starts at time t=0 and moves along a number line so that its position, at time t greater than or equal to 0, is given by x(t)=(t-2)^3(t-6). The particle is moving to the right for 1) 0 < t < 5 2) 2 < t < 6 3) t >5 4) t greater than or equal to 0 5) Never

OpenStudy (michele_laino):

here we have to solve this inequality: \[x\left( t \right) > 0 \Rightarrow {\left( {t - 2} \right)^3}\left( {t - 6} \right) > 0\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

i did that and i got (-infinity to 2) U (6, infinity)

OpenStudy (michele_laino):

yes! correct! More precisely, since \(t \geqslant 0 \), we have: \[0 \leqslant t < 2 \cup t > 6\]

OpenStudy (michele_laino):

nevertheless it is not an option of yours

OpenStudy (anonymous):

So the answer will be 5) because there is no option like that

OpenStudy (michele_laino):

I think taht maybe the right function \(x(t)\), can be this: \[x\left( t \right) = {\left( {t - 2} \right)^3}\left( {6 - t} \right)\]

OpenStudy (anonymous):

thank you this was more to just check my answer i wanted to make sure i understand what to do in a situation like this one

OpenStudy (michele_laino):

that*

OpenStudy (anonymous):

I understand your concern however the function given was x(t)= (t-2)^3(t-6)

OpenStudy (michele_laino):

in that case I don't know how to think, since we have found the time interval, wherein the particle is moving to the right

OpenStudy (michele_laino):

try to ask to your teacher, then let me know his reply :)

OpenStudy (anonymous):

I will try my best to do that Sir

OpenStudy (michele_laino):

ok! :)